the coffee shoppe sells a coffee blend made from two coffees, one costing $5/lbs and the other costing $6/lbs. If the blended coffee sells for $5.60/lb find how much of each coffee is used to obtain the desired blend. Assume that the weight of the blended coffee is 100lbs.
Answer by nerdybill(7384)

the coffee shoppe sells a coffee blend made from two coffees, one costing $5/lbs and the other costing $6/lbs. If the blended coffee sells for $5.60/lb find how much of each coffee is used to obtain the desired blend. Assume that the weight of the blended coffee is 100lbs.
.
Let x = amount (lbs) of $5 coffee
then
100-x = amount (lbs) of $6 coffee
then
5x + 6(100-x) = 5.60(100)
5x + 600-6x = 560
600-x = 560
-x = -40
x = 40 lbs ($5 coffee)
.
$6 coffee:
100-x = 100-40 = 60 lbs